If the mean and the stadard deviation of a continuous random variable that is normally distributed are 50 and 4, respectively, find an interval that contains 58% of the distribution.

Answer by khwang(438)

Assume the 95% interval is (mu -x,mu+x) where mu is the mean = 50
We have x/4 = 0.806 [ this value corresponding to 0.58 from the Table]
so, x = 0.806*4 = 3.26
Hence,the solution interval is (50- 3.26,50+3.26)
= (46.77, 53.26)
Kenny